OpenStudy (anonymous):

Identify the vertex for the graph of y = -3x2 + 6x - 2.

OpenStudy (anonymous):

is it 1,1

OpenStudy (mertsj):

x coordinate of vertex is -b/2a=-6/-6=1 Replace x with 1 to find the y coordinate of the vertex. -3(1)^2+6(1)-2=-3+6-2=1 Vertex is (1,1)